OpenStudy (vortish):

use the quadratic formula to find the x-intercpets on the the graph of the equation y=x^2+8x-5

OpenStudy (anonymous):

For the equation of the form ax^2 + bx + c = 0, the x-intercepts will be \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (vortish):

it say math processing error dr v

OpenStudy (tkhunny):

Do you have a = 1, b = 8 and c = -5?

OpenStudy (vortish):

so it would be 8/2(1)

OpenStudy (vortish):

Then it woud be x=4

OpenStudy (vortish):

any thoughts

OpenStudy (tkhunny):

?? What did you do to get x = 4? With a = 1, b = 8 and c = -5, we have: \(x = \dfrac{-8\pm\sqrt{8^{2}-4(1)(-5)}}{2(1)}\)

OpenStudy (vortish):